OpenStudy (firejay5):

Simplify. Show work and explain. 12. a + b over a^2 - b^2

OpenStudy (mertsj):

What are the factors of 25-v^2

OpenStudy (firejay5):

I think I figured it out. I think I know the answer

OpenStudy (mertsj):

ok good for you.what did you get?

OpenStudy (firejay5):

- v + 5 over 3v + 2

OpenStudy (mertsj):

\[\frac{(5-v)(5+v)}{(3v+2)(v-5)}=\frac{-(5+v)}{3v+2}\]

OpenStudy (firejay5):

didn't I have that

OpenStudy (mertsj):

Missing a parentheses

OpenStudy (mertsj):

-(5+v) is not the same as -v+5

OpenStudy (firejay5):

\[-\frac{ v + 5 }{ 3v + 2 }\]

OpenStudy (mertsj):

yep

OpenStudy (firejay5):

that's what I had typed in before, except it was hard to read

OpenStudy (firejay5):

@Mertsj changed question

OpenStudy (mertsj):

Good job!!

OpenStudy (firejay5):

it's 2/3 <---- 4ay over 6ay ---> cancel ay from numerator and denominator and you get 4/6; which reduces to 2/3

OpenStudy (mertsj):

yes

OpenStudy (firejay5):

sorry changed question again

OpenStudy (mertsj):

np

OpenStudy (mertsj):

What did you get?

OpenStudy (firejay5):

1 over a - b

OpenStudy (mertsj):

You're rollin' now

OpenStudy (firejay5):

am I really?

OpenStudy (mertsj):

Looks like it to me

OpenStudy (firejay5):

really

OpenStudy (firejay5):

I use Wolram|Alpha to check/get the answer and then I show my work

OpenStudy (mertsj):

Very good.

OpenStudy (firejay5):

am I good or what

OpenStudy (mertsj):

no doubt

OpenStudy (firejay5):

what's that suppose to mean

OpenStudy (mertsj):

No doubt you are good!!